Assuming that 30% of all industrial accidents in a certain plant are caused by failure of employees to follow instructions, find, approximately, the probability that among 84 industrial accidents in this plant anywhere from 20 to 30 (inclusive) will be due to failure of employees to follow instructions.
Answer by Boreal(15235)

mean is np=25.2
variance is np(1-p)=25.2*0.70=17.64
sd = sqrt (V)=4.2
z(min)=(19.5-25.2)/4.2
=-5.7/4.2=-1.36
-
z=(30.5-25.2)/4.2
=5/3/4.2
=1.26
probability z between those two values is 0.8093